Use the information given below to calculate the net wage of H, Kane for the last week of May 2021.


INFORMATION:

H, Kane is employed by Hot- Spur Enterprises. The normal working week is 45 hours.

H, Kane worked for 48 hours during the last week of May 2021.

She is remunerated at R80 per hour during normal working hours.

Overtime is calculated at 1½ times the normal rate.


The following are her deductions for the week:

  • Medical aid fund R350
  • Pension fund 8% of the normal wages
  • Income tax R830

Hot- Spur Enterprises contributes the same amount as the employees towards the medical aid fund and pension fund.


1
Expert's answer
2022-03-27T07:24:30-0400

Normal wage=45x80

=R3600

Overtime wage= 3x1.5x80

=R360

Total weekly wage= 3600+360

=R3960


Deductions:

Medical= R350

pension=8/100x3600=R288

income tax=R830

Total Deductions=R1568

Net wages=3960-1568

=R2392